# Deeper quantum circuits via pseudo-twirling coherent errors mitigation in non-Clifford gates

### v1

@@ V Summary and concluding remarks
The working hypothesis underlying the present paper is that shorter gates facilitate deeper circuits. However, the coherent errors in short multi-qubit gates cannot be addressed with randomized compiling (Pauli twirling) techniques as it hinges on the assumption that the ideal gates are Clifford gates. Our analytical, numerical, and experimental findings indicate that pseudo twirling can bridge this gap and successfully mitigate coherent errors in non-Clifford gates. Pseudo twirling possesses three advantages that separate it from the established randomized compiling approach: i) Applicability to Non-Clifford Gates: It enables the implementation of significantly shorter gates, resulting in lower noise accumulation compared to Clifford-based implementations. ii) Intra-Gate Coherent Error Suppression: Pseudo twirling allows for the suppression of coherent errors within gates and sliced pulses, leading to substantial improvement in coherent error suppression. iii) Circuit-PST: By applying pseudo twirling also at the edges of a circuit, it is possible to suppress non-local multi-qubit coherent errors.
Moreover, the successful integration of PST with the Adaptive-KIK error mitigation method leads to a substantial performance boost. This boost can be instrumental in achieving quantum advantage in a wide range of fields, including particle physics, condensed matter, quantum chemistry, and more.
@@ Acknowledgements.
The authors acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. We thank IAI/ELTA Systems for the assistance in running the experiments. Raam Uzdin is grateful for support from the Israel Science Foundation (Grant No. 2556/20).
@@
where $\rho^{\prime}$ is the transformed density matrix and $\{K_{i}\}$ are Kraus operators that satisfying the relation $\sum_{i}K_{i}^{\dagger}K_{i}=I$ (being $I$ the $d\times d$ identity matrix). In the Hilbert space formalism, observables are described by a hermitian operators $A$. The expectation value of $A$ for system in the state $\rho$ reads
@@
The Liouville space representation serves as an alternative mathematical formulation that greatly aids in streamlining the expression of quantum states and their transformations. In this model, the density matrix is transformed into a state vector, indicated as $|\rho\rangle$, with a dimensionality of $1\times d^{2}$, and a quantum transformation is depicted by a superoperator matrix with dimensions $d^{2}\times d^{2}$.
utilize the scripted notation $\mathcal{O}$ for a generic quantum operation, its corresponding expression in Liouville space, Eq. (3), is replaced by
@@
Following the methodology of Ref. [30], $|\rho\rangle$ is the column vector whose first $d$ components correspond to the first row of $\rho$, the next $d$ components correspond to the second row of $\rho$, and so forth. More explicitly the “vec-ing” procedure a generic $d\times d$ matrix $B$ reads $|B\rangle=\left(B_{11},B_{12},...,B_{1d},B_{21},B_{22},...,B_{2d},...,B_{d1},B_{d2},...,B_{dd}\right)^{\textrm{T}}$, where $B_{ij}$ is the $ij$ entry of $B$. With this convention, in Liouville space a Kraus map (3) takes the form [30]
@@
where $K_{i}^{\ast}$ is the element-wise complex conjugate of $K_{i}$. Since unitary map is a special case of Kraus map $\rho^{\prime}=U\rho U^{\dagger}$, is Liouville space a unitary evolution is given by
$|\rho^{\prime}\rangle=\mathcal{U}|\rho\rangle=U\otimes U^{\ast}|\rho\rangle$ (7)
Equation (6) is follows from the vectorization rule for a product of three matrices $B$, $C$ and $D$ [30]
$|BCD\rangle=B\otimes D^{\textrm{T}}|C\rangle,$ (8)
@@
Lastly, expectation value (4) of an observable matrix $A_{d\times d}$ can be neatly expressed in Liouville space using the ’bra’ of $\left\langle A\right|=\left|A\right\rangle^{\dagger}$. That is, as a column row vector $\langle A|=\left(A_{11}^{\ast},A_{12}^{\ast},...,A_{1d}^{\ast},...,A_{d1}^{\ast},A_{d2}^{\ast},...,A_{dd}^{\ast}\right)$. As result, using the hermiticity of $A$ we obtain that of the expectation value of $A$ can be written as a ’bra-ket’ inner product of $\left|A\right\rangle$ and $\left|\rho\right\rangle$
@@
Appendix II - The issue with using virtual gates in PST
@@
$Ue^{-\frac{i}{2}\theta\sigma_{z}}\to e^{-\frac{i}{2}\theta\sigma_{z}}e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}=e^{-\frac{i}{2}\theta\sigma_{z}}[e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}]$
For simplicity, we assume that $U=e^{+i\phi P_{\alpha}}$ where $P_{\alpha}$ is some Pauli. If $[P_{\alpha},\sigma_{z}]=0$, $Ue^{-i\theta\sigma_{z}}=e^{-i\theta\sigma_{z}}U$, then we are done since we have pushed the $e^{-i\theta\sigma_{z}}$ to the next layer. The idea in virtual gate is to keep pushing the $R_{z}$ gates all the way to the end of the circuit. Since the measurement is in the computational basis, the final $R_{z}$ rotation has no impact.
@@
$=\cos^{2}\frac{\theta}{2}P_{H}+i\sin\frac{\theta}{2}[\sigma_{z},P]+\sin^{2}\frac{\theta}{2}\sigma_{z}P\sigma_{z}.$
@@
$(\cos^{2}\frac{\theta}{2}-\sin^{2}\frac{\theta}{2})P+i2\sin\frac{\theta}{2}\cos\frac{\theta}{2}\sigma_{z}P)=\cos\theta P+i\sin\theta\sigma_{z}P.$
@@
$e^{+\frac{i}{2}\theta\sigma_{z}}\sigma_{x}e^{-\frac{i}{2}\theta\sigma_{z}}=\cos\theta\sigma_{x}+i\sin\theta\sigma_{z}\sigma_{x}=\cos\theta\sigma_{x}+\sin\theta\sigma_{y},$
which means that the physical implementation of $e^{+\frac{i}{2}\theta\sigma_{z}}Pe^{-\frac{i}{2}\theta\sigma_{z}}$ just involves changing the phase of the pulse generating $P_{\alpha}$. For example, an $R_{z}$ before the controlled qubit of ZX cross resonance pulse, will manifest in changing the phase of the CR pulse.
Let us consider the PST case
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}$
In pseudo twirling, we use Pauli gates so $\theta=\pi$, and therefore the virtualization will amount to changing the sign of the drive without affecting the coherent error
$\mathcal{P}_{IZ_{v}}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}.$
@@
$\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}=e^{+\mathcal{P}_{IZ}(i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh})\mathcal{P}_{IZ}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{P}_{IZ}\mathcal{H}_{coh}\mathcal{P}_{IZ}},$
@@
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{YI}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{YI}.$
However, if $\mathcal{P}_{YI}$ is generated via virtual Z: $\mathcal{P}_{IY}=\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}$ we obtain
@@
$\mathcal{P}_{IX}e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IX}.$
Thus, the IY twirling is indistinguishable from IX twirling. While aiming to sample from the full set of twirlings, the use of virtual Z and also Y using virtual Z, decimates the effective set of accessible twirlings and consequently degrades the ability of PST to mitigate the full scope of coherent errors.

### v2

@@ VI Summary and concluding remarks
The working hypothesis underlying the present paper is that shorter gates facilitate deeper circuits. However, the coherent errors in short multi-qubit gates cannot be addressed with randomized compiling (Pauli twirling) techniques as it hinges on the assumption that the ideal gates are Clifford gates. Our analytical, numerical, and experimental findings indicate that pseudo twirling can bridge this gap and successfully mitigate coherent errors in non-Clifford gates. Pseudo twirling possesses three advantages that separate it from the established randomized compiling approach: i) Applicability to non-Clifford gates: it enables the implementation of significantly shorter gates, resulting in lower noise accumulation compared to Clifford-based implementations. ii) Intra-gate coherent error suppression: pseudo twirling allows for the suppression of coherent errors within gates and sliced pulses, leading to substantial improvement in coherent error mitigation. iii) Edges-PST: by applying pseudo twirling also at the edges of a circuit, it is possible to suppress non-local multi-qubit coherent errors. Moreover, the successful integration of PST with the Adaptive-KIK error mitigation method leads to a substantial performance boost. This boost can be instrumental in achieving quantum advantage in a wide range of fields, including particle physics, condensed matter, quantum chemistry, and more.
@@ Acknowledgements.
The authors acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. We thank IAI/ELTA Systems for the assistance in running the experiments. Raam Uzdin is grateful for support from the Israel Science Foundation (Grant No. 2556/20). The support of the Israel Innovation Authority is greatly appreciated.
@@
where $\rho^{\prime}$ is the transformed density matrix and $\{K_{i}\}$ are Kraus operators that satisfying the relation $\sum_{i}K_{i}^{\dagger}K_{i}=I$ ($I$ being the $d\times d$ identity matrix). In the Hilbert space formalism, observables are described by a Hermitian operators $A$. The expectation value of $A$ for system in the state $\rho$ reads
@@
The Liouville space representation serves as an alternative mathematical formulation that greatly aids in streamlining the expression of quantum states and their transformations. In this model, the density matrix is transformed into a state vector, indicated as $|\rho\rangle$, with a dimensionality of $d^{2}\times 1$, and a generic quantum transformations $\mathcal{O}$ is described by a superoperator matrix with dimensions $d^{2}\times d^{2}$ such that
@@
Following the methodology of Ref. [40], $|\rho\rangle$ is the column vector whose first $d$ components correspond to the first row of $\rho$, the next $d$ components correspond to the second row of $\rho$, and so forth. More explicitly, the vectorization procedure of a $d\times d$ matrix $B$ reads $|B\rangle=\left(B_{11},B_{12},...,B_{1d},B_{21},B_{22},...,B_{2d},...,B_{d1},B_{d2},...,B_{dd}\right)^{\intercal}$, where $B_{ij}$ is the $ij$ entry of $B$. With this convention, in Liouville space a Kraus map (28) takes the form [40]
@@
where $K_{i}^{\ast}$ is the element-wise complex conjugate of $K_{i}$. Since unitary maps $\rho^{\prime}=U\rho U^{\dagger}$ are a special case of Kraus map , in Liouville space a unitary evolution is given by
$|\rho^{\prime}\rangle=\mathcal{U}|\rho\rangle=U\otimes U^{\ast}|\rho\rangle.$ (32)
Equation (31) follows from the vectorization rule for a product of three matrices $B$, $C$ and $D$ [40]
$|BCD\rangle=B\otimes D^{\intercal}|C\rangle,$ (33)
@@
Lastly, expectation value (29) of an observable matrix $A_{d\times d}$ can be neatly expressed in Liouville space using the ’bra’ vector $\left\langle A\right|=\left|A\right\rangle^{\dagger}$. That is, as a row vector $\langle A|=\left(A_{11}^{\ast},A_{12}^{\ast},...,A_{1d}^{\ast},...,A_{d1}^{\ast},A_{d2}^{\ast},...,A_{dd}^{\ast}\right)$. Using the Hermiticity of $A$ we obtain that of the expectation value of $A$ can be written as a ’bra-ket’ inner product of $\left|A\right\rangle$ and $\left|\rho\right\rangle$
@@
Appendix II - Derivation of the relation $\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}=0$
[…]
$\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}$ $=\sum_{\alpha\in\beta^{+}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}+\sum_{\alpha\in\beta^{-}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha},$ (35)
where $\beta^{\pm}$ is the set of $\alpha$’s such that: $\mathcal{P}{}_{\alpha}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha}=\pm\mathcal{H}_{\beta}$ ($sgn(\alpha,\beta)=\pm 1$). Further simplification leas to
$\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}$ $=\sum_{\alpha\in\beta^{+}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}-\sum_{\alpha\in\beta^{-}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}$
$=2\sum_{\alpha\in\beta^{+}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}.$ (36)
The last equality follows from the twirling property $\sum_{\alpha}\mathcal{P}{}_{\alpha}\mathcal{H}\mathcal{P}{}_{\alpha}=0$ that holds for any $\mathcal{H}$. Next, we define a two-element permutation within $\beta^{+}$. To each element $\mathcal{P}{}_{\alpha_{1}^{+}}$ in $\beta^{+}$ we can assign another Pauli operator $\mathcal{P}{}_{\alpha_{2}^{+}}$ in $\beta^{+}$ using the transformation
$\mathcal{P}{}_{\alpha_{2}^{+}}=\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}},$ (37)
where $\mathcal{P}{}_{\delta}$ is chosen such $\mathcal{P}{}_{\delta}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}=\mathcal{H}_{\beta}$ and $\mathcal{P}{}_{\delta}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}=-\mathcal{H}_{\gamma}$, i.e. $[P_{\delta},P_{\beta}]=0$ and $\{P_{\delta},P_{\gamma}\}=0$. To verify that $\alpha_{2}^{+}\in\beta^{+}$ we check if $\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha_{2}^{+}}=+\mathcal{H}_{\beta}$
$\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{P}{}_{\delta}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha_{1}^{+}}=+\mathcal{P}{}_{\alpha_{1}^{+}}.$ (38)
which confirms that $\alpha_{2}^{+}\in\beta^{+}$. In the derivation we use the fact that in Liouville space all $\mathcal{P}$ commute with each other. We proceed by studying the outcome of $\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{2}^{+}}$ compared to $\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{1}^{+}}$
$\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{2}^{+}}=\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{P}{}_{\delta}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=-\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{1}^{+}}.$ (39)
We conclude that every element $\alpha\in\beta^{+}$ has counterpart element that has opposite sign to the sum (36). As a result, the sum is zero. In the more general case $g_{\gamma}\mathcal{H}_{\gamma}\to\sum_{\gamma}g_{\gamma}\mathcal{H}_{\gamma}$ a different $P_{\delta}$ can be chosen to each element in the sum so the proof above applies automatically to this case as well. Finally, we point out that if it happens that $\{P_{\beta},P_{\gamma}\}=0$ it is possible to choose $P_{\delta}=P_{\beta}$.
Appendix III - derivation of $\mathcal{K}_{pst}^{(m)}$
Under the assumption that the slices are sufficiently short that neither the drive or coherent error change substantially along the slice we can write
$\mathcal{K}_{pst}^{(m)}=e^{-idt\mathcal{H}(t_{1})}e^{+\mathcal{L}_{dt}^{pst}}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}.....e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}},$ (40)
where $dt=T/m$, $t_{m}=T$, and $\mathcal{L}_{dt}^{pst}=dt^{2}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})$ (see Eq. (14)). By inserting the identities $\{U_{k}U_{k}^{\dagger}\}_{k=1}^{m}$ and $U_{m-k}^{\dagger}e^{-idt\mathcal{H}(t_{k+1})}=U_{m-k-1}^{\dagger}$ we obtain:
$\mathcal{K}_{pst}$ $=U_{m}[U_{m}^{\dagger}e^{-idt\mathcal{H}(t_{1})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-1}][U_{m-1}^{\dagger}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-2}]$
$\times$ $[U_{m-2}^{\dagger}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-3}][U_{m-3}^{\dagger}.....e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}}$
$=U_{m}[U_{m-1}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-1}][U_{m-2}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-2}][U_{m-3}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-3}]$
$\times$ $U_{m-3}^{\dagger}.....U_{1}U_{1}^{\dagger}e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}}$
$=U_{m}e^{U_{m-1}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-1}^{\dagger}}e^{U_{m-2}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-2}^{\dagger}}e^{U_{m-1}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-1}^{\dagger}}...e^{+\mathcal{L}_{dt}^{pst}}.$ (41)
By keeping the leading order in $\mathcal{L}_{dt}^{pst}$ we have
$\mathcal{K}_{pst}=\mathcal{U}_{T}e^{\sum_{k}\mathcal{U}_{m-k}^{\dagger}\mathcal{L}_{dt}^{pst}\mathcal{U}_{m-k}}.$ (42)
In regular dissipation $\mathcal{L}_{dt}\propto dt$ so in the limit of large m we obtain the standard first Magnus term. However the present case $\mathcal{L}_{dt}^{PST}\propto dt^{2}=\frac{T}{m}dt$ and therefore additional $T/m$ coefficient appears and we have
$\mathcal{K}_{pst}^{(m)}$ $=\mathcal{U}_{T}e^{\frac{T}{m}\sum_{k=0}^{m-1}\mathcal{U}_{k}^{\dagger}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}_{k}dt}$
$\to\mathcal{U}_{T}e^{\frac{T}{m}\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}(t)dt}.$ (43)
Appendix IV - using the operator norm for quantifying errors
In our simulations we use the operator norm to describe the error in the evolution operator. In this appendix we show that it is directly related to the error in measuring observables. The operator norm of a matrix $S$ is equal to its largest singular value. Yet, an alternative definition is:
$\left\|S\right\|_{op}=max_{\psi}\frac{|\left\langle\psi\left|S^{\dagger}S\right|\psi\right\rangle|}{\sqrt{\left\langle\psi\left|\psi\right.\right\rangle}}.$ (44)
Applying the Cauchy–Schwarz inequality to the error in some expectation value $\langle A|\rho_{f}\rangle=\langle A|\mathcal{K}|\rho_{0}\rangle$ with respect to the ideal value $\langle A|\mathcal{U}|\rho_{0}\rangle$ we have:
$\left|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle\right|$ $\leq\sqrt{\left\langle A\left|A\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|(\mathcal{K}-\mathcal{U})^{\dagger}(\mathcal{K}-\mathcal{U})\right|\rho_{0}\right\rangle}$
$\leq\sqrt{\left\langle A\left|A\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}\left\|\mathcal{K}-\mathcal{U}\right\|_{op},$ (45)
where we used (44) in the last step. Since the constants $\sqrt{\left\langle A\left|A\right.\right\rangle}$ and $\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}$ are trivial to evaluate, we conclude that $\left\|\mathcal{K}-\mathcal{U}\right\|_{op}$ sets a clear and easy to calculate bound on the error in any observable $A$ and any initial state $\rho_{0}$. Notably, often the diamond norm is used for comparing different maps. While the diamond norm has some appealing features that can help in proofs, it is difficult to evaluate it when the number of qubits increases. In contrast, the operator norm is trivial to calculate. Furthermore, in the example discussed in Fig. 5, we found that the operator norm and the diamond norm deviate by less than ten percent without any features that distinct one from each other.
Finally, we point out that (45) can be slightly refined when $\mathcal{K}$ is a trace-preserving map. For trace preserving maps $\left\langle I\left|\mathcal{U}\right|\rho_{0}\right\rangle=\left\langle I\left|\mathcal{K}\right|\rho_{0}\right\rangle=1$ and therefore $|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle|=|\left\langle A+\alpha I\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle|$. This implies that $\sqrt{\left\langle A\left|A\right.\right\rangle}$ can be replaced by $\sqrt{\left\langle A+\alpha I\left|A+\alpha I\right.\right\rangle}$. Minimizing over $\alpha$ we find that
$\left|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle\right|$ $\leq\sqrt{\left\langle A^{\prime}\left|A^{\prime}\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}\left\|\mathcal{K}-\mathcal{U}\right\|_{op},$ (46)
$A^{\prime}$ $=A-tr(A)I/tr(I).$ (47)
Appendix V - the issue with using virtual gates in PST
@@
$Ue^{-\frac{i}{2}\theta\sigma_{z}}\to e^{-\frac{i}{2}\theta\sigma_{z}}e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}=e^{-\frac{i}{2}\theta\sigma_{z}}[e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}].$ (48)
For simplicity, we assume that $U=e^{+i\phi P_{\alpha}}$ where $P_{\alpha}$ is a Pauli Matrix. If $[P_{\alpha},\sigma_{z}]=0$, $Ue^{-i\theta\sigma_{z}}=e^{-i\theta\sigma_{z}}U$ and the $e^{-i\theta\sigma_{z}}$ was pushed to the next layer. The idea in virtual gate is to keep pushing the $R_{z}$ gates all the way to the end of the circuit. Since the measurement is in the computational basis, the final $R_{z}$ rotation has no impact.
@@
$=\cos^{2}\frac{\theta}{2}P+i\sin\frac{\theta}{2}[\sigma_{z},P]+\sin^{2}\frac{\theta}{2}\sigma_{z}P\sigma_{z}.$ (49)
@@
$(\cos^{2}\frac{\theta}{2}-\sin^{2}\frac{\theta}{2})P+i2\sin\frac{\theta}{2}\cos\frac{\theta}{2}\sigma_{z}P)=\cos\theta P+i\sin\theta\sigma_{z}P.$ (50)
@@
$e^{+\frac{i}{2}\theta\sigma_{z}}\sigma_{x}e^{-\frac{i}{2}\theta\sigma_{z}}=\cos\theta\sigma_{x}+i\sin\theta\sigma_{z}\sigma_{x}=\cos\theta\sigma_{x}+\sin\theta\sigma_{y},$ (51)
which means that the physical implementation of $e^{+\frac{i}{2}\theta\sigma_{z}}Pe^{-\frac{i}{2}\theta\sigma_{z}}$ just involves changing the phase of the pulse generating $P_{\alpha}$. For example, an $R_{z}$ before the target qubit of ZX cross-resonance pulse, will manifest in changing the phase of the cross-resonance pulse.
Let us now consider how virtual Z work in PST
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}.$ (52)
In pseudo twirling, we use Pauli gates so $\theta=\pi$ in $R_{z}(\theta)$, and therefore the virtualization will amount to changing the sign of the drive without affecting the coherent error
$\mathcal{P}_{IZ_{v}}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}.$ (53)
@@
$\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}=e^{+\mathcal{P}_{IZ}(i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh})\mathcal{P}_{IZ}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{P}_{IZ}\mathcal{H}_{coh}\mathcal{P}_{IZ}},$ (54)
@@
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{IY}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IY}.$ (55)
However, if $\mathcal{P}_{IY}$ is generated via virtual Z: $\mathcal{P}_{IY}=\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}$ we obtain
@@
$\mathcal{P}_{IX}e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IX}.$ (56)
Thus, the IY twirling is indistinguishable from IX twirling. While aiming to sample from the full set of twirlings, the use of virtual Z and also Y using virtual Z, decimates the effective set of accessible twirlings and consequently degrades the ability of PST to mitigate the full scope of uncontrolled coherent errors.